I'm at a loss with this one because I am not sure if I should distribute first and then factor after moving the terms to the left to = 0 or not.
Solve. (x^2 - 2x)(x + 3) = -2x(x+1)

Answer by CharlesG2(834)

I'm at a loss with this one because I am not sure if I should distribute first and then factor after moving the terms to the left to = 0 or not.
Solve. (x^2 - 2x)(x + 3) = -2x(x+1)
use FOIL on left, use distribution on right
x^3 + 3x^2 - 2x^2 - 6x = -2x^2 - 2x
now arrange and group like terms
x^3 + 3x^2 - 2x^2 + 2x^2 - 6x + 2x = 0
x^3 + 3x^2 - 4x = 0 (factor out the x)
x(x^2 + 3x - 4) = 0
x(x + 4)(x - 1) = 0 (by FOIL (x+4)*(x-1) does equal above)
x = 0 or x = -4 or x = 1